OpenStudy (anonymous):

Solve for the missing side in the following right triangle: hypotenuse = 17, leg = 8, leg = ?. A. 13 B. 15 C. 12 D. 14

OpenStudy (ivettef365):

use pythagorean theorem

OpenStudy (ivettef365):

17^2 = 8^2 + b^2, solve for b

OpenStudy (anonymous):

usePythagorean theorem

OpenStudy (anonymous):

b^2=17^2-8^2 b^2=289-64